I have asked this question before but may not have written it properly. Can someone please help?
Solve the matrix equation:
|4-5||m|=|32|
|1 2||n|=|-5|

Answer by stanbon(48568)

4-5||m|=|32|
|1 2||n|=|-5|

I wonder what method you are studying.
The coefficient determinant is D=(8--5)=13
The m-determinant is Dm=(64-25)=39
The n-determinat is Dn=(-20-32)=-52
Then m= Dm/D=39/13 = 3
And n=Dn/D = -52/13 = -4
Cheers,
Stan H.